a ball is thrown forward horizontally from the top of a cliff with a velocity of 10 ms^-1 the height of the cliff above the ground is 45m calculate:

1. the time to reach the ground

2. the distance from the cliff of the ball on hitting the ground

3. the direction of the ball to the horizontal just before it hits the ground


1
Expert's answer
2021-10-10T15:58:28-0400
"t=\sqrt{2\frac{45}{9.8}}=3.0\ s\\d=(10)(3)=30\ m\\
\theta=\arctan{\frac{9.8\cdot3}{10}}=71\degree\text{ with horizontal}"